# LATTE: Latent Trajectory Embedding for Diffusion-Generated Image Detection

### v1

@@ 4.1 Experimental Setup
Datasets and pre-processing. We conduct our experiments on the GenImage dataset [19]  [^1] Licensed under CC BY-NC-SA 4.0., which comprises real and fake images generated by 8 different generative models: BigGAN [20], GLIDE [4], VQDM [52], Stable Diffusion V1.4 and V1.5 [7], ADM [53], Midjourney [54], and Wukong [55]. While GenImage serves as our primary benchmark, we also include the Diffusion Forensics [13]  [^2] License not specified. dataset for additional cross-generator evaluation. All images are resized to $224\times 224$ before being passed to the model.
Evaluation metrics. Following standard practice in generated image detection, we evaluate the effectiveness of our models using accuracy (ACC) and average precision (AP). These metrics provide a comprehensive view of both overall classification performance and precision-recall trade-offs.
Training details. We extract the latent trajectories using Stable Diffusion 2.1333Available under the CreativeML Open RAIL-M license: https://huggingface.co/stabilityai/stable-diffusion-2-1. as a denoising model. We empirically chose five evenly spaced timesteps: [981, 741, 521, 261, 1] for extracting the latents. The visual features are encoded with a pretrained ConvNeXt [56]  [^4] Available via the timm library under the Apache 2.0 license: https://huggingface.co/timm/convnext_base.fb_in22k. yielding a dimension size of $512$. All models are trained by minimizing binary cross-entropy loss to convergence, monitored on a held-out validation split matching the training generator. We use a batch size of 32, AdamW [57] optimizer (learning rate = 1e-4, weight decay = 4e-5), and a cosine-annealed learning rate scheduler. The experiments are conducted on a single H100 GPU, by training approximately 2 hours for a single epoch. The code is available at: https://github.com/AnaMVasilcoiu/LATTE-Diffusion-Detector.
@@ 4.2 Comparison to Baselines
We first evaluate our proposed method on GenImage [19], which is essentially testing cross-generator generalization performance. Specifically, we train our model on a given subset and test on all other subsets, and then we repeat this across all eight generators. Table 1 presents the averaged accuracy of our method in comparison to a range of existing baselines, while Figure 3 provides a complete breakdown of the results, in terms of accuracy and average precision, across all subsets. It can be observed that LATTE achieves the highest average accuracy, exceeding or matching the baselines. In particular, on one of the most challenging subsets, such as BigGAN — which is the only GAN-based generator — LATTE outperforms LaRE by $13.8$%. This demonstrates that explicitly modeling trajectory evolution in the latent space yields more generalizable detection across diverse generators.
Table 1: Accuracy(%) comparison of LATTE and other baselines on GenImage. Each value represents the averaged accuracy across the test sets of each generator. LATTE attains top-tier average accuracy, demonstrating the potential of the latent trajectory modeling.
@@
CNNSpot [58] 58.2 70.3 70.2 57.0 57.1 67.7 56.7 56.5 61.7
F3Net [59] 55.1 73.1 73.1 66.5 57.8 72.3 62.1 56.5 64.6
Spec [60] 56.7 72.4 72.3 57.9 65.4 70.3 61.7 64.3 65.1
ResNet-50 [61] 59.0 72.3 72.4 59.7 73.1 71.4 60.9 66.6 66.9
DIRE [13] 65.0 73.7 73.7 61.9 69.1 74.3 63.4 56.7 67.2
DeiT-S [62] 60.7 74.2 74.2 59.5 71.1 73.1 61.7 66.3 67.6
Swin-T [63] 61.7 76.0 76.1 61.3 76.9 75.1 65.8 69.5 70.3
LaRE [14] 66.4 87.3 87.1 66.7 81.3 85.5 84.4 74.0 79.1
@@
Figure 3: Accuracy(%) and average precision(%) of LATTE across the GenImage dataset. The x-axis indicates the generator used to produce the training data, while each bar represents the model’s performance when tested on data from the different generators.
4.3 Cross-Domain Generalization
Additionally, we evaluate the cross-domain generalization ability of LATTE, and compare it to our baseline LaRE [14]. We are testing the best-performing model trained on the SDv1.4 subset of GenImage against the Diffusion Forensics dataset, introduced in [13]. For a fair comparison, we retrain LaRE from scratch using the same training data as LATTE. Table 2c reports the accuracies across various generators and three distinct domains - Bedroom, ImageNet, and CelebA - which differ substantially from GenImage in both content and style, offering a robust benchmark for evaluating transferability.
Table 2: Cross-domain generalization results of models trained on GenImage and tested on Diffusion Forensics. LATTE achieves an overall average improvement of 7.1% over LaRE.
@@ 4.3 Cross-Domain Generalization
Our model, LATTE, consistently outperforms LaRE across all three domains, achieving improvements of 16.2% on Bedroom, 4% on Imagenet and 1.1% on Celeba, with an overall average gain of 7.1%.
@@ 4.4 Ablation Study
Influence of selected timesteps.
@@
Table 3: Accuracy(%) comparison of varying latent trajectory resolutions. We compare the effect of different timestep configurations (Tn) on the average classification accuracy across eight generative models. The best accuracy is achieved with the 5-timestep setting (T5).
Timesteps Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
T1 80.4 99.2 98.8 71.2 85.6 96.9 74.2 62.5 83.6
T3 78.7 99.3 99.0 75.3 81.8 97.2 80.4 68.0 85.0
@@
T9 86.2 99.5 99.4 75.2 94.4 98.0 79.8 86.9 89.9
Influence of vision backbone choice. We examine the impact of the vision backbone on performance. Standard CLIP encoders (RN50, ViT-B/32) underperformed on GenImage in preliminary experiments, prompting the shift to higher-capacity backbones: ConvNeXt-Base pretrained on ImageNet-22k and CLIP ViT-L/14  [^5] Available under the MIT License: https://github.com/openai/CLIP., the latter inspired by UnivFD [64]. Both improved the results significantly, with ConvNeXt consistently achieving the highest accuracy, as demonstrated in Table 4.
Table 4: Accuracy(%) comparison between different vision backbones. ConvNeXt outperforms CLIP ViT-L/14 by 5.2%.
@@
Benefit of average pooling strategy. The standard average pooling used to summarize timestep-specific representations assumes equal importance across all timesteps in the trajectory. To challenge this assumption, we experiment with a learned weighted pooling mechanisms that assigns importance scores to each timestep using a linear gating function and softmax normalization. While this allows different timesteps to have varying influence on the final representation, it performs worse than simple average pooling - suggesting that, in practice, all steps contribute equally informative signals.
@@
In addition, we consider aggregation through CLS pooling. Namely, we replace the average pooling with a self-attention over the sequence of refined latents with added positional encodings. We prepend a [CLS] token that serves as the trajectory aggregation. The goal is to assess whether allowing latents to refine each other via self-attention and incorporate sequence order can improve performance. However, this variant underperforms compared to simple average pooling, suggesting that the refined LATTE is already expressive enough without additional self-attention refinements, as seen in Table 5.
@@
Table 5: Accuracy(%) comparison between different aggregation strategies. Average pooling outperforms learnable weighted pooling by 9.3% and CLS pooling by 1.3%.
@@
4.5 Embedding Space Analysis
Besides quantitative results, we assess our model’s discriminative capacity by visualizing real and generated image embeddings with t-SNE [65]. The first row of Figure 4 depicts the embeddings extracted from the original, frozen ConvNeXt backbone, while the second row displays embeddings after the backbone has been fine-tuned during LATTE training.
Figure 4: Visualizations of t-SNE embeddings for real and fake images across five generators from GenImage. The first row presents embeddings before using LATTE (extracted using the original ConvNeXt), while the second row shows embeddings derived from LATTE. The reduced overlap in the second row indicates LATTE’s discriminative power and generalization capabilities.
The embeddings in the second row exhibit much clearer separation between real (blue) and fake (orange) samples, indicating reduced overlap and stronger class separation. This clear separation demonstrates that LATTE’s joint training captures the latent trajectory dynamics and adapts the visual encoder to better distinguish generated content.
4.6 Robustness to Perturbations
Additionally, we assess the robustness of LATTE under common post-processing operations like compression, resizing, Gaussian blur, and Gaussian noise. Such perturbations often occur in real-world pipelines and can severely degrade the subtle artifacts that detection methods depend on. As shown in Figure 5, LATTE consistently outperforms LaRE, maintaining higher detection accuracy and greater stability. This shows that LATTE’s reliance on multi-step latent trajectories is more invariant under such transformations than single-step reconstruction errors.
Figure 5: Accuracy(%) of LATTE and LaRE on perturbed images. We evaluate and compare the robustness of both methods under four common post-processing transformations: JPEG compression, Center crop & resize, Gaussian blur, and noise. LATTE consistently outperforms LaRE across all considered perturbations.
4.7 Qualitative Analysis
To complement the quantitative results, we present a selection of qualitative examples in Figure 6, organized in a confusion-matrix-style layout. This visualization provides insight into typical model successes and failure cases, illustrating the visual characteristics that influence detection outcomes and highlighting both the strengths and limitations of the proposed approach.
Top-left: True negatives. This quadrant displays real images that the model correctly identifies as real. These examples contain complex textures, human subjects, or high-frequency structural details. Such content remains inherently difficult for generative models to replicate with full fidelity, resulting in strong detection cues. This observation reflects a broader trend across the GenImage dataset, namely that real-world images characterized by organic structure, intricate details, or recognizable human elements are consistently classified correctly, highlighting the model’s strength in identifying authentic content.
Top-right: False Positives. These are real images that the model mistakenly classifies as fake. A common pattern among these examples is the presence of smooth textures, highly saturated colors, or stylized lighting, visual characteristics often associated with generative outputs. Real images that stylistically resemble synthetic content, such as those with artistic post-processing, exaggerated lighting, or minimal texture complexity, tend to trigger false alarms. This reflects a known limitation, where the model may overestimate the likelihood of manipulation in real images that exhibit generative-like visual properties.
Bottom-left: False Negatives. This quadrant contains generated images that the model fails to detect, incorrectly classifying them as real. Many of these examples depict simple objects, clean compositions, or scenes lacking complex structure or obvious artifacts. When generated content exhibits high visual plausibility and avoids common generative flaws, the model’s detection performance degrades. Such borderline cases, where fakes are indistinguishable from real content, particularly when lacking intricate textures or irregularities, remain a challenge for our latent trajectory-based detection method. These examples underscore the difficulty of reliably identifying high-quality synthetic images and highlight the growing need for robust, adaptive detectors, especially in light of the rapid advancements in generative model capabilities.
Bottom-right: True Positives. These are generated images correctly identified as fake by the model. Despite appearing visually convincing at first glance, the subtle generative traces embedded in the latent denoising trajectory leveraged by LATTE enable reliable detection. Our detector does not rely on superficial visual artifacts but instead exploits inconsistencies inherent to the generative process, often imperceptible to humans. While high-quality synthetic images remain a significant challenge for detection methods, as highlighted in the previous quadrant, these examples illustrate that LATTE takes an important step toward overcoming this difficulty, enabling reliable identification even for visually plausible, high-fidelity fakes.
5 Conclusion
We introduced LATTE, the first approach to explicitly model the trajectory of latent embeddings across diffusion denoising steps. Our two-stage architecture - (i) Latent-Visual Feature Refinement and (ii) Latent-Visual Fusion & Classification - yields performance gains over prior work and improved generalization capabilities. In particular, LATTE achieves an average accuracy of 82.2% on the GenImage dataset (a 3.1% improvement over LaRE), and delivers a 7.1% gain in cross-domain accuracy on Diffusion Forensics, underscoring the effectiveness of leveraging latent trajectory dynamics for robust generated-image detection. We believe these findings pave the way for future research into latent trajectory-aware detection features for generalizable generated image detection.
Limitations. While LATTE achieves strong performance and improved generalization, its effectiveness diminishes under strong post-processing (e.g., heavy JPEG compression or strong blur), indicating sensitivity to distribution shifts. Additionally, like most global detectors, LATTE has been evaluated primarily on fully synthetic versus real images. Detecting small, localized forgeries is a different task and remains an open challenge for future work.
Broader Impact. Our work contributes to the detection of diffusion-generated images, supporting efforts to counter disinformation and enhance media authenticity. However, as with any detection method, it may prompt adversarial adaptation or be misused in contexts where real content is unjustly flagged. Future work should consider robust evaluation in real-world situations and mitigation strategies for misuse.
Supplementary material
Appendix A Complete in-domain and cross-model comparison to the baselines
In this section, we report a comprehensive matrix of pairwise detection results to complement the summary metrics in the main paper. For DIRE [13] and LaRE [14], we copy the results reported in the LaRE paper. For our method, LATTE, we evaluate it under the same experimental protocol to enable fair comparison. In each case, the detector is trained on images generated by one model and tested on held-out images from eight generators, producing an $8\times 8$ matrix. Diagonal entries reflect in-domain performance (i.e. training and testing on the same generator), while off-diagonal entries measure cross-model generalization.
Figure 7: Complete pairwise evaluation of detection performance across all 8 generators in the GenImage dataset. Each subplot corresponds to one detector - DIRE (left; baseline), LaRE (center; baseline), and LATTE (right; proposed) - and shows the accuracy(%) when training on the subset listed on the vertical axis and testing on the subset listed along the horizontal axis. Row- and column-averages summarize each method’s cross-model generalization capabilities.
DIRE achieves a strong in-domain average accuracy of 94.24%, but it’s cross-model generalization is limited, with an off-diagonal mean of only 63.5%. LaRE substantially improves both metrics - reaching 99.96% in-domain and 76.08% off-diagonal. Our method, LATTE, maintains competitive in-domain performance at 98.89% while further boosting transfer - with an average cross-model accuracy of 80.18%, outperforming LaRE by nearly 4 percentage points. This demonstrates that incorporating temporal trajectory consistency not only preserves high fidelity on known generators but also yields the most robust detection on unseen models.
Appendix B Additional ablation studies
To further understand the key design choices and components of the LATTE framework, we conduct a series of additional ablation studies. These experiments are aimed at isolating and quantifying the contributions of individual architectural modules, processing strategies, and training configurations. All ablation results reported in this section are based on models trained using the SDv1.4 subset of the GenImage dataset. We focus on this subset since using a fixed generator allows us to isolate the effects of individual design choices without confounding variability, and SDv1.4 is a commonly used baseline in prior work (e.g. [16]), making our results easier to contextualize.
B.1 Importance of each component in LATTE
We perform an ablation study focusing on three key elements: the visual feature $\mathbf{v}_{\text{IMG}}$ from the backbone, the latent trajectory derived from intermediate diffusion steps, and the Latent-Visual Feature Refinement module that aligns and integrates these modalities via cross-attention.
We evaluate four model variants with different combinations of these components. The first baseline (A) uses visual features only, while the second (B) uses only the latent trajectory, aggregated with mean pooling. The third (C) combines the visual features with the latent trajectory, allowing the model to capture generative dynamics, but without any refinement. The final variant (D) includes all three components.
Table 6: Ablation on visual and latent components. ✓ indicates the component is included. Results are shown as Accuracy (%) for each generator. Including all components outperform the visual and latent-only configurations by 16% and 37.7%.
Model Visual Latent Refine MJ SDV1.4 SDV1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
A 83.5 99.9 99.9 51.7 56.2 99.9 58.4 50.0 74.9
B 50.0 58.3 58.4 50.0 50.0 56.6 52.6 50.0 53.2
C 80.5 100 99.9 76.7 84.3 99.8 69.2 75.8 85.7
D 88.7 99.8 99.7 74.0 95.7 98.9 80.8 89.7 90.9
As shown in Table 6, we observe that both visual-only (A) and latent-only (B) models yield relatively poor performance, indicating that neither visual semantics nor latent diffusion signals alone are sufficient for robust detection. Adding latent trajectory features to the visual backbone (C) leads to improvements across most generators, suggesting that these two modalities offer complementary cues. However, the accuracy gains are still limited, likely due to misalignment between the latent and visual representations when no refinement is applied. The full model (D), which incorporates both latent trajectories and the Latent-Visual Feature Refinement module, achieves the highest accuracy across nearly all subsets. Notably it outperforms other variants by a large margin on challenging generators (e.g. +11.6% on VQDM, +13.9% on BigGAN). These results highlight the critical role of refinement in effectively fusing latent and visual information.
B.2 Influence of separate latent processing strategy
Table 7: Accuracy(%) comparison between separate vs. joint latent processing strategies. Processing timesteps separately yields the highest average accuracy, outperforming joint processing by 0.7%.
Joint 88.4 99.7 99.6 72.5 94.4 98.6 79.4 88.9 90.2

### v2

@@ 4.1 Experimental Setup
Datasets. We evaluate LATTE across several complementary settings. First, to assess overall detection and cross-generator robustness, we use the GenImage dataset [^1] Licensed under CC BY-NC-SA 4.0.(Zhu et al., 2023b), which contains real and fake images generated by eight generative models, including diffusion- and GAN-based approaches. Next, to test performance under more visually-challenging scenarios, we use the Chameleon dataset (Yan et al., 2025), which includes high-quality synthetic images designed to reduce detection artifacts. To evaluate cross-domain generalization, we use the Diffusion Forensics dataset (Wang et al., 2023), which spans multiple visual domains such as bedrooms, churches, and faces. All images are resized to $224\times 224$ before being passed to the model.
Training & Evaluation. We extract the latent trajectories using Stable Diffusion 2.1. We empirically chose five timesteps: [981, 741, 521, 261, 1] for extracting the latents, which evenly spread across the trajectory. The visual features are obtained with a pretrained ConvNeXt (Liu et al., 2022), yielding a dimension size of $512$. All models are trained by minimizing binary cross-entropy loss to convergence, monitored on a held-out validation split matching the training generator. We use a batch size of 32, AdamW (Loshchilov & Hutter, 2017) optimizer (lr = 1e-4, weight decay = 4e-5), and a cosine-annealed learning rate scheduler. The experiments are conducted on a single H100 GPU, by training for approximately 2 hours for a single epoch. To provide a comprehensive evaluation, we follow standard practice in detection tasks and evaluate our models using accuracy and average precision. The code repository, training, and evaluation details will be released.
@@ 4.2 Comparison to Baselines
We first evaluate our method on GenImage (Zhu et al., 2023b), which essentially tests cross-generator generalization. All models are trained on images generated by SDv1.4 and evaluated across eight different generators, with baseline results cited from Yan et al. (2025). The results, shown in Table 1, indicate that LATTE/Avg (using average pooling) achieves the highest average accuracy among a broad set of related methods, improving by 4.1% over the recent AIDE model, followed by LATTE/CLS (using CLS token pooling). Notably, on one of the most challenging subsets - BigGAN, LATTE/Avg surpasses the strongest prior baseline (Ojha et al., 2023) by 9.59.5%. Note that we continue using LATTE/Avg in the subsequent experiments as our best model, denoted as LATTE for brevity.
Table 1: Comparison of LATTE to baselines on GenImage benchmark (Zhu et al., 2023b). All methods are trained on SDv1.4 of GenImage and evaluated over eight image generators. LATTE/Avg achieves the best average accuracy, improving by $4.1\%$ over state-of-the-art methods.
@@
CNNSpot (Wang et al., 2020) 52.8 96.3 95.9 50.1 39.8 78.6 53.4 46.8 64.2
F3Net (Qian et al., 2020) 50.1 99.9 99.9 49.9 50.0 99.9 49.9 49.9 68.7
Spec (Zhang et al., 2019) 52.0 99.4 99.2 49.7 49.8 94.8 55.6 49.8 68.8
GramNet (Liu et al., 2020) 54.2 99.2 99.1 50.3 54.6 98.9 50.8 51.7 69.9
DeiT-S (Touvron et al., 2021) 55.6 99.9 99.8 49.8 58.1 98.9 56.9 53.5 71.6
ResNet-50 (He et al., 2016) 54.9 99.9 99.7 53.5 61.9 98.2 56.6 52.0 72.1
DIRE (Wang et al., 2023) 65.8 99.7 99.7 54.5 58.1 99.4 54.3 49.8 72.7
UnivFD (Ojha et al., 2023) 73.2 84.2 84.0 55.2 76.9 75.6 56.9 80.3 73.3
Swin-T (Liu et al., 2021) 62.1 99.9 99.8 49.8 67.6 99.1 62.3 57.6 74.8
GenDet (Zhu et al., 2023a) 89.6 96.1 96.1 58.0 78.4 92.8 66.5 75.0 81.6
DRCT (Chen et al., 2024) 94.6 99.8 99.8 61.8 65.9 99.9 74.8 58.8 82.1
PatchCraft (Zhong et al., 2023) 79.0 89.5 89.3 77.3 78.4 89.3 83.7 72.4 82.3
Co-Spy (Cheng et al., 2025) 83.4 96.8 96.7 67.2 93.0 95.9 78.8 65.2 84.6
LaRE (Luo et al., 2024) 74.0 100 99.9 61.7 88.5 100 97.2 68.7 86.2
AIDE (Yan et al., 2025) 79.4 99.7 99.8 78.5 91.8 98.7 80.3 66.9 86.9
LATTE/CLS 85.8 99.0 99.0 83.2 86.7 96.8 88.7 77.8 89.6
@@
Next, we train our model on each generator-specific subset of GenImage and test it across all other subsets. Figure 3 reports the averaged accuracies, where our model again achieves the best overall performance. These findings demonstrate that explicitly modeling the trajectory evolution in latent space yields stronger robustness and more reliable detection across diverse image generators.
Figure 3: Comparison of LATTE to baselines, by training and testing across all 8 generators of GenImage. Each plot corresponds to one detector - DIRE (left; baseline), LaRE (center; baseline), and LATTE (right; proposed) - and shows the accuracy(%) when training on the subset listed on the vertical axis and testing on the subset listed along the horizontal axis.
We further evaluate our model on Chameleon (Yan et al., 2025), a recently proposed benchmark designed to reflect real-world scenarios by covering a broad range of content, including humans, animals, objects, and scenes. This benchmark allows us to test how well the model generalizes beyond its training distribution and captures transferable representations. As shown in Table 2, our method achieves consistent improvements over the baselines, achieving 2.5% improvement over AIDE when trained on the GenImage dataset. The results highlight both the robustness of our approach and its effectiveness in generalizing across diverse visual domains.
Finally, we evaluate the cross-domain generalization ability of LATTE on the Diffusion Forensics dataset (Wang et al., 2023). Specifically, we train LATTE on the SDv1.4 subset of GenImage and use LaRE and AIDE trained on the same data for a fair comparison. Table 3c reports accuracies across various generators and three distinct domains - Bedroom, ImageNet, and CelebA - which differ substantially from GenImage in both content and style. Across all three domains, LATTE consistently outperforms both LaRE and AIDE, achieving improvements such as 11.1% on Bedroom and 4% on Imagenet, with an overall average gain of 7.1%. We also notice that in-domain performance (train and test on the same data) is already saturated in prior work - often reported at or near 100% - so it offers limited insight into generalization. Therefore, we prioritized the evaluation in cross-generator and cross-domain settings.
Table 2: Cross-domain comparison on Chameleon (Yan et al., 2025). Each column represents the accuracy (%) of different detectors, and the rows indicate the used training set.
[…]
Table 3: Cross-domain comparison on Diffusion Forensics (Wang et al., 2023). LATTE achieves an overall average improvement of 7.1% accuracy over LaRE and 14.8% over AIDE.
@@ 4.3 Ablation Study
In this section, we present ablation studies to quantify the contribution of each component, the impact of the denoising steps and vision backbone. Additional ablations are available in the Appendix A.
Importance of each component. We conduct an ablation study on three components: the visual features from the backbone, the latent trajectory from intermediate diffusion steps, and the Latent–Visual Fusion module that aligns these modalities via cross-attention. Four model variants are evaluated: (A) visual features only, (B) latent trajectory only, (C) visual + latent trajectory without fusion, and (D) the full model with all components. As shown in Table 4, both visual-only (A) and latent-only (B) variants perform poorly, confirming that neither modality alone is sufficient. Combining the two in (C) improves performance, indicating complementary cues, but the gains remain limited. The full model (D) achieves the best results across nearly all subsets, with large improvements on challenging cases such as VQDM (+11.6%) and BigGAN (+13.9%), underscoring the importance of effectively fusing latent and visual information.
Table 4: Ablation on visual and latent components. ✓ indicates that the component is included. Results are shown as Accuracy (%) for each generator. Including all components of our approach outperforms the visual-only and latent-only configurations by 16.1% and 37.8%.
Model Visual Latent Fusion Midjourney SDV1.4 SDV1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
A 83.5 99.9 99.9 51.7 56.2 99.9 58.4 50.0 74.9
B 50.0 58.3 58.4 50.0 50.0 56.6 52.6 50.0 53.2
C 80.5 100 99.9 76.7 84.3 99.8 69.2 75.8 85.7
@@
Influence of denoising steps.
We study how performance changes with the number of denoising steps, varying $n\in\{1,3,5,9,13,15\}$ used to sample intermediate latents for the trajectory. For $n=5$ steps, we empirically select the following: [981, 741, 521, 261, 1], while $n=1$ corresponds to the single midpoint t = 521. The remaining configurations include both endpoints ($t=1$ and $t\approx 1000$) with additional steps interpolated evenly across the trajectory. Our choice of such evenly spaced steps - spanning from near the start to the end of the trajectory - aims to capture the full spectrum of denoising behavior. As shown in Table 5, accuracy improves as the number of sampled steps increases, peaking at $n=5$. Beyond this point, the $n=9$ configuration maintains competitive results, but performance declines at $13$ and $15$ steps, suggesting that adding more steps introduces redundancy rather than additional useful information.
Table 5: Accuracy(%) comparison of varying lengths of latent trajectory. We compare the effect of different timestep configurations on the average accuracy across eight generative models. The best accuracy is achieved with the 5-timestep configuration ($n=5$).
$n$-steps Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
1 80.4 99.2 98.8 71.2 85.6 96.9 74.2 62.5 83.6
3 78.7 99.3 99.0 75.3 81.8 97.2 80.4 68.0 85.0
@@
9 86.2 99.5 99.4 75.2 94.4 98.0 79.8 86.9 89.9
13 77.3 99.7 99.4 72.2 82.8 98.5 77.3 69.4 84.6
15 77.3 99.6 99.3 73.5 82.9 98.4 78.3 62.7 84.0
Influence of vision backbone. In our preliminary experiments, we used CLIP encoders (RN50, ViT-B/32), which underperformed on GenImage. This prompted the shift to other backbones: ConvNeXt-Base (Liu et al., 2022) pretrained on ImageNet-22k and CLIP ViT-L/14 (Ilharco et al., 2021), also leveraged by Ojha et al. (2023). Both improved the results significantly, with ConvNeXt consistently achieving the highest accuracy, as demonstrated in Table 6.
Table 6: Accuracy(%) comparison between different vision backbones. ConvNeXt outperforms CLIP ViT-L/14 by 5.3%.
@@
4.4 Embedding Space Analysis
We assess our model’s discriminative capacity by visualizing real and generated image embeddings with t-SNE (van der Maaten & Hinton, 2008) in Figure 4. The first row depicts the embeddings extracted from the original, frozen ConvNeXt backbone, while the second row displays embeddings after the backbone has been fine-tuned with LATTE. The embeddings in the second row exhibit much clearer separation between real (blue) and generated (orange) samples, indicating reduced overlap and stronger class separation.
Figure 4: Visualizations of t-SNE embeddings for real and fake images across five generators from GenImage. The first row presents embeddings before using LATTE (extracted using the ConvNeXt), while the second row shows embeddings derived from LATTE. The much clearer separation in the second row illustrates LATTE’s discriminative power.
4.5 Robustness to Unseen Perturbations
We assess the robustness of LATTE under common post-processing operations like compression, resizing, Gaussian blur, and Gaussian noise. Such perturbations often occur in real-world pipelines and can severely degrade the subtle artifacts that detection methods depend on. As shown in Figure 5, LATTE consistently outperforms LaRE, maintaining higher detection accuracy and greater stability. This shows that LATTE’s reliance on multi-step latent trajectories is more invariant under such transformations than single-step reconstruction errors.
Figure 5: Accuracy(%) of LATTE vs. LaRE on perturbed images. We evaluate and compare the robustness of both methods under four common transformations: JPEG compression, center crop & resize, Gaussian blur, and noise. LATTE consistently outperforms LaRE across all perturbations.
4.6 Qualitative Analysis
We present qualitative examples in a confusion-matrix-style layout in Figure 6, highlighting representative model successes and failures. Top-left: Real images with complex textures, human subjects, or fine structures are typically recognized as authentic, since such details remain difficult for generative models to replicate. Top-right: In contrast, some real images with smooth textures, saturated colors, or stylized lighting are misclassified as fake, reflecting the model’s sensitivity when authentic content visually resembles synthetic imagery. Bottom-left: On the other hand, high-quality generated images that appear simple or artifact-free may be mistaken for real, highlighting the difficulty of detecting visually convincing fakes. Bottom-right: Lastly, LATTE succeeds in correctly identifying visually convincing fake images, which suggests that it leverages subtle traces rather than only visual artifacts.
5 Conclusion
We propose LATTE, a novel diffusion-generated image detection approach that models the sequential evolution of latents across multiple denoising steps. By capturing trajectory patterns and grounding them with visual features, LATTE learns a compact and discriminative representation. Experiments on GenImage, Chameleon, and Diffusion Forensics demonstrate that LATTE achieves state-of-the-art performance, including significant gains in cross-generator and cross-domain scenarios. Overall, this work highlights latent trajectory modeling as a new direction for generated image detection.
Limitations. While LATTE achieves strong performance and improved generalization, its effectiveness diminishes under strong post-processing (e.g., heavy JPEG compression or strong blur), indicating sensitivity to distribution shifts. Additionally, like most global detectors, LATTE has been evaluated primarily on fully synthetic versus real images, while detecting small, localized forgeries remains a distinct challenge for future work.
Ethics Statement
This work advances the field of synthetic media forensics by improving the detection of generated images. As generative models improve their ability to produce highly realistic content, frameworks or tools like LATTE, play an important role in combating disinformation, verifying content authenticity, and maintaining public trust in digital media.
However, the deployment of the detection system also raises important ethical and societal considerations. As detection technologies improve, so do adversaries’ strategies for evading them, potentially resulting in an arms race between generation and detection. Furthermore, there is a risk that such tools will be misapplied, for example, by incorrectly labeling legitimate content as false or by being employed in politically or socially biased ways. Overreliance on automated systems is another growing concern, as they may miss edge cases or fail silently in unfamiliar situations.
Reproducibility Statement
We are committed to ensuring the reproducibility of our results. To this end, we will release the full source code and evaluation scripts upon publication. Our paper clearly and fully describes the proposed feature extraction method and the model architecture in Section 3, and provides comprehensive details on the experimental setup in Section 4.1, including used datasets, preprocessing steps, training configurations, and hyperparameters. Additionally, ablations and variant evaluations in Section 4.3 and Appendix A further support reproducibility.
Appendix
The appendix consists of the following sections: A. Additional Ablation Studies, B. Latent Trajectory Spatial Analysis, C. Complete Accuracy and AP on GenImage, D. Architectural Details of the CLS-pooling, and E. Embedding Space Analysis.
Appendix A Additional ablation studies
To further understand the key design choices and components of the LATTE framework, we conduct a series of additional ablation studies. All ablation results reported in this section are based on models trained using the SDv1.4 subset of GenImage (Zhu et al., 2023b).
A.1 Benefit of average pooling
@@
Standard pooling in LATTE assumes equal importance across all timesteps in the trajectory. To test this design choice, we experiment with a weighted pooling mechanism that assigns importance scores to each timestep using a linear gating function and softmax normalization. As shown in Table 7, this approach performs worse than simple average pooling - suggesting that all steps provide equally informative signals. We also consider CLS pooling, where a special token aggregates the sequence of latents through self-attention with positional encodings. The goal is to assess whether allowing the latents to refine each other via self-attention and incorporating sequence order can improve performance. This variant slightly underperforms, suggesting that LATTE is already expressive enough without additional attention-based aggregations.
@@
Table 7: Accuracy(%) comparison between different aggregation strategies. Average pooling outperforms learnable weighted pooling by 9.4% and CLS pooling by 1.4%.
@@
A.2 Effect of latent extraction configuration
The sequence of latents is obtained by first encoding real and fake images into latent space using a frozen VAE, followed by partial reconstruction via a pre-trained diffusion model. At each timestep, noise is added to the VAE latents and then partially denoised via the UNet, capturing intermediate latent representations along the reconstruction trajectory.
We ablate two factors in this latent extraction pipeline: the choice of sampling method (DDPM vs. DDIM, Table 8) and the choice of U-Net backbone (Stable Diffusion v1.5 vs. v2.1, Table 9). For the sampling method, we use Stable Diffusion v2.1 as the backbone, while for the U-Net model comparison, we fix the scheduler to DDPM.
Table 8: Accuracy(%) comparison between DDPM and DDIM-based latent extraction. DDPM improved accuracy by 7.2%.
Sampling method Midjourney SDv1.4 SDv1.5 ADM GLIDE Wukong VQDM BigGAN Avg.
DDIM 77.0 99.7 99.5 74.6 82.2 98.2 77.4 61.8 83.8